OpenStudy (anonymous):

solve ^3sqrt2x+4+9=1

OpenStudy (anonymous):

please type correctly...

OpenStudy (anonymous):



OpenStudy (anonymous):

the 9 is not under the thing

OpenStudy (anonymous):

-258

OpenStudy (anonymous):

\[\sqrt[3]{2x+4}\] = 1-9 \[\sqrt[3]{2x+4}\] = -8 \[\sqrt[3]{2x+4}\] = -1 * 8 \[\sqrt[3]{2x+4}\] = -1 * \[\sqrt[3]{8}\] 2x+4 = -1 * 2 2x + 4 = -2 2x = -6 x = -3

OpenStudy (anonymous):

How can cube root of 2x+4 be equal to -1*8 and -1*cube root of 8?

OpenStudy (anonymous):

-8 is'nt cube root

OpenStudy (anonymous):

If x is substituted as -3, then the answer doesn't result to 1. Check your calculations again.

OpenStudy (anonymous):

then show ur steps

OpenStudy (anonymous):